### v1

# Efficient Transformer-based Hyper-parameter Optimization for Resource-constrained IoT Environments

Abstract
The hyper-parameter optimization (HPO) process is imperative for finding the best-performing Convolutional Neural Networks (CNNs). The automation process of HPO is characterized by its sizable computational footprint and its lack of transparency; both important factors in a resource-constrained Internet of Things (IoT) environment. In this paper, we address these problems by proposing a novel approach that combines transformer architecture and actor-critic Reinforcement Learning (RL) model, TRL-HPO, equipped with multi-headed attention that enables parallelization and progressive generation of layers. These assumptions are founded empirically by evaluating TRL-HPO on the MNIST dataset and comparing it with state-of-the-art approaches that build CNN models from scratch. The results show that TRL-HPO outperforms the classification results of these approaches by 6.8% within the same time frame, demonstrating the efficiency of TRL-HPO for the HPO process. The analysis of the results identifies the main culprit for performance degradation attributed to stacking fully connected layers. This paper identifies new avenues for improving RL-based HPO processes in resource-constrained environments.
Index Terms:
Hyper-parameter Optimization, Reinforcement Learning, Transformers, Image Classification, MNIST dataset, Resource constraints, IoT Environment
I Introduction
Convolutional Neural Networks (CNNs) are the staple implementation of neural networks (NNs) for image classification and object detection. The progress in this field is attributed to the increased complexity of CNN architectures in terms of the type of connections among different layers and their depth, which increased the computational demands. As a result, the accuracy of these models depends on the number and the type of layers, the connections between these layers, and the parameters assigned to each layer [1]. In addition to accuracy, other important factors that dictate the design of CNNs include training time, inference time, and energy consumption, which are emphasized in the resource-constrained Internet of Things (IoT) environment. The number of possible choices for designing CNNs is extremely large, which is prohibitive for manual search. These circumstances have promoted the automated CNN architecture search, achieved through the field of Neural Architecture Search (NAS).
The NAS field has recently seen great progress, due to the incorporation of Reinforcement Learning (RL) agents to search for the best CNN configurations. The RL’s appeal stems from its generalizability of hyper-parameter (HP) combinations via function approximation and the trial-and-error approach, which has the potential to reduce the computational demand of NAS [2]. This field is dominated by the literature that either builds CNNs from scratch using basic NN layers or a pre-defined collection of these layers, or optimizes the HPs of an already-existing CNN.
Autonomous vehicles (AV) are part of the envisioned IoT applications that utilize edge servers of limited computational resources. The realization of AVs heavily depends on image classification models deployed on these servers. As such, the disparate computational capabilities of edge servers are prohibitive for the prolonged execution of large models. Model partitioning is key to resolving these limitations toward fulfilling the promise of AVs [3]. This requires gathering insights into the contribution of each CNN model’s layers to the classification results. In current implementations, this vision is stalled by multiple oversights. The sequential nature of model generation results from the dependence between layer combinations. This dependence can be reflected by the recurrent structures of Recurrent Neural Networks (RNNs). Therefore, RNN-based controllers constructing CNN models dominated the RL-based NAS field [4, 5]. The one-at-a-time processing of inputs inhibits their parallelization, resulting in their insurmountable computational footprint that is ill-fitted for the resource-scarce vehicular environment. On a different note, these controllers are built to receive a reward when a terminal condition is encountered, masking the contribution of each layer to prediction results. Within the constraints of computational resources and model partitioning requirements, the state-of-the-art (SOTA) approaches are limited by their impracticability and lack of transparency, hindering their deployment in real-world environments.
This paper presents Transformer-based Reinforcement Learning Hyper-parameter Optimization (TRL-HPO) that alleviates the shortcomings of the current methods. The TRL-HPO addresses the transparency and long-computational times of RL-based solutions with competitive performances. The Transformer architecture overshadows the RNN-based methods by integrating the multi-headed self-attention mechanism [6]. The TRL-HPO controller equipped with the attention mechanism allows us to gain insights from the layer-generation process, enhancing its interpretability. Moreover, the transformers’ architecture supports parallelization, which addresses the long computation times of its RNN-based counterparts [6]. The reward is produced with every generated layer to showcase improvement instead of waiting for termination conditions to unfold, a condition matching CNN partitioning requirements, as depicted in Figure 1. The contributions of this paper are as follows:
  • Propose a novel Hyper-parameter Optimization (HPO) process named TRL-HPO that is the first to combine transformer architecture and actor-critic RL;
  • Enhance the CNN’s model performance generated within a shorter period compared to SOTA approaches using the TRL-HPO process;
  • Improve the transparency of the CNN model generation process by examining the attention-reward affinities and their layer combinations;
  • Create open challenges related to RL-based HPO processes, including TRL-HPO, for researchers to address.
The rest of the paper is organized as follows. Section II presents the related work and its limitations. Section III details the proposed framework. Section IV explains the implementation details and evaluation criteria. Section V analyzes the obtained results. Section VI concludes the paper.
II Related Work
The field of hyper-parameter optimization (HPO) is an important research topic in Machine Learning (ML) with practical implementations in real-world environments that enhance the performance of Deep Neural Networks (DNNs).
Bayesian Optimization (BO), Evolutionary Search (ES) algorithms, and RL agents are the three main tools for HPO implementation. BO and ES methods are limited by their assumptions [1] and lack of generality [7], which favors the application of RL techniques. The works of Baker et al. [5] and Zoph et al. [4] are the first to propose incorporating RL methods into generating CNN models. The former work utilized RNN-based controllers using an off-policy RL algorithm to sample CNN architectures while the latter utilized a value-based Q-learning approach. The results of these seminal works highlighted the trade-off between the running time of HPO methods and the accuracy of the obtained models.
The computational footprint of RL implementations for HPO promoted the utilization of multi-agent RL that shrinks the state space of each agent. The work of Neary [8] uses multiple agents to optimize HPs of the CNNs built from scratch, whereby a master agent orchestrates the synchronization between the other agents outputting HPs. On the other hand, the work in [9] assigns each agent to optimize the HPs of an already-existing CNN layer, such that the dependence in HP space is mapped using a shared Q-table between consecutive layers. While all of these studies focus on a single objective, the works presented by Hsu et al. [10] and Tan et al. [11] incorporate multiple objectives in the reward function formulation. MONAS [10] considers the energy and accuracy constraints, whereas MNASNet [11] integrates the inference latency of developed models.
The shortcomings exhibited by the integration of RL into the HPO process in these implementations hinder their deployment in IoT environments. First, the adopted models lack the transparency that shows the interdependence between different layers. Second, some of these approaches take considerable time to converge, which is a privilege that is not always available in resource-constrained environments. Third, some works focus on optimizing the HPs of a specific layer, reducing the state space at the expense of the layer’s diversity. Lastly, many of the works include prior knowledge instilled into the layer-generation process, such as the addition of dropout layers [4, 5]. These limitations are addressed in this paper by proposing a transformer-based RL controller and the reward function formulation. The multi-headed attention mechanism facilitates the training process and adds transparency to the model generation process by analyzing the attention values. On the other hand, the reward function reflects the contribution of each layer to enhancing the classification results, a condition favorable to CNN model partitioning.
[…]
[scale=0.45]architecture_v3.PNG
Figure 1: TRL-HPO framework
III Methodology: TRL-HPO
This section explains the building blocks of TRL-HPO, including the transformer and RL’s actor-critic architectures, which are shown in Figure 1, and the motivation for this combination toward a more efficient and transparent HPO process.
III-A Transformer
The vanilla transformer is a sequence-to-sequence model, which avoids the recurrence structures of RNNs by integrating the innovative self-attention mechanism [6]. This architecture enables parallelization accelerating the training of transformer models. The building blocks of transformers include attention mechanism, multi-headed attention, position-wise feed-forward network, and positional encoding (PE). The transformers follow the encoder-decoder structure thereby each of these structures is formed by stacking combinations of these identical layers. The ordering of these layers with respect to the encoder and decoder is depicted in Figure 1.
The Positional Encoding computes the sequence’s order and is added to the input embedding of the encoder and decoder stacks so that the order is incorporated into the input and output data. To obtain a unique order, a sinusoidal function takes as inputs the position and the embedding dimensions. This structure extrapolates to sequences that are longer than the ones encountered during training [6]. The self-attention module calculates the weights assigned by each input sequence to other sequences. This way the recurrence structure is avoided because self-attention reflects the dependencies between input and specific output and improves the modelling of long-range dependencies [6]. As such, the attention mechanism resembles a fully connected layer (FCL), such that the weights reflect the pairwise relationship from previous inputs. To fully exploit this method, the transformer includes multi-headed attention that replicates the attention calculation. This process generates different representations of the input data at different positions. The calculation of these attention matrices is independent and facilitates their parallelization. Lastly, a feed-forward neural network is used as the subsequent layers to the multi-head attention mechanism.
III-B Actor-Critic Reinforcement Learning
RL teaches an agent to perform a task by accumulating experiences from interacting with its environment. Its actions are refined based on a reward function that signals the utility of these actions [12]. In connection to the HPO process, the RL’s experimentation with the absence of any prior knowledge mirrors the inability to hypothesize the best combination of basic CNN layers. The only knowledge for the HPO process relates to the order of the stacked layer that starts with grid-like inputs and ends with a fully connected layer (FCL). As such, the combination of unknown orientation and trial-and-error experimentation matches the requirements of the HPO process.
The RL methods can be split into value-based and policy-based methods. The value-based method estimates the quality of a state-action pair using a value function and optimizes this value iteratively using actions that maximize it. The high variability of value estimations and their under-performance in scenarios with continuous action spaces promotes policy-based methods [12]. The policy-based methods focus on optimizing the policy to maximize the expected cumulative rewards [12]. However, the latter methods suffer from sample inefficiency, which is an asset of value-based methods. To complement these methods’ advantages, an actor-critic approach combines both policy-based (actor) and value-based (critic) methods [12]. Projected to the HPO problem, the actor outputs an action that maps to a layer and its HPs. The critic evaluates this action by outputting a value showing the quality of the taken action [12].
III-C Transformers and Actor-critic RL
As the name suggests, the proposed framework, Transformer-based Reinforcement Learning HPO (TRL-HPO) is the convergence of transformers and actor-critic RL. The actor and the critic are implemented using the transformer architecture. The steps involved in the actions’ generation, CNN model construction, and their evaluation are depicted in Figure 1. The combination of TRL-HPO enables the RL agent to harness the strengths of transformers to the benefit of the HPO process. This new framework is the first trial to integrate transformers into the HPO process, opening a new frontier toward the exploitation of this novel architecture to address a lingering problem in the field of ML. This experiment does not plan to only exhibit the viability of transformers’ implementation to HPO, but it rather analyzes the suitability of this architecture to the HPO process.
We analyze TRL-HPO based on its inherent advantages and its benefits versus the SOTA approaches. In terms of the framework, two main advantages are reaped. The first is that TRL-HPO builds models from scratch, rendering it a general-purpose method. The proof of concept is constrained to a single use case; however, the framework can be expanded to any problem that utilizes DNN models. The second concerns the definition of the reward function that is progressively updated with each stacked layer. While this definition incurs an extra computational footprint by training progressively deeper models, it facilitates understanding the contribution of each layer. This feature fits the requirements of environments that seek to balance resource constraints and accuracy objectives. The transformer architecture results in two inherent advantages to the HPO process. The first is connected to the parallelization enabled by the multi-headed attention mechanism, which reduces the running time of the HPO process. The second benefit stems from the attention mechanism that showcases the relationships between generated layers. A by-product of this feature is the inference of the combination of layers that produce better results, which enhances the transparency of the HPO process.
IV Implementation Details
The evaluation of TRL-HPO is conducted on the MNIST dataset [13]. This dataset is a large database of handwritten digits, from 0 to 9, containing 60,000 training images and 10,000 testing images. Each image is a greyscale image of 28 $\times$ 28 dimensions. This dataset serves as a good proof-of-concept for the proposed framework, due to its availability and limited complexity. The former factor enables a fair comparison with other SOTA approaches, such as [5, 8]. The latter factor facilitates training CNN models in a limited time; an advantage when working with hardware constraints. The comparison with SOTA approaches is based on the convergence time to the best solution (hrs), the classification accuracy (%), and the best CNN classification model (%) obtained by each SOTA approach upon the completion of TRL-HPO process (AccTime). We compare TRL-HPO with Baker et al. [5] and Neary [8], which are the only RL-based solutions that build CNN models from scratch.
Layer HP Values
CNN filters {8, 16, …, 128}
kernel {3, 5, 7}
stride {1, 2, 3}
FCL neurons {16, 24, …, 512}
bias {T, F}
activations
{None, relu, leakyrelu,
tanh, sigmoid, elu, gelu}
MaxPool kernel {2, 3, …,, 8}
stride {1, 2, 3}
padding {0, 1, 2, 3}
TABLE I: Set of hyper-parameters
The generated CNN models are constrained to 6 layers, formed as a combination of the three basic layers: CNN layer, FCL layer, and Maximum Pooling (MaxPool) layer. The set of HPs used in the experimental procedure is summarized in table I. Stacking a combination of these layers adapts to the input structure. In connection to the RL agent, the design of the state and action space and the reward function is imperative to drive the RL agents’ action generation, which requires defining the transformers’ input sequence.
The reward function represents a layer’s contribution towards improving the classification results, which means that the DNN model’s first layer produces the highest rewards compared to any subsequent layers. Each layer is represented as a combination of the representation of the layer itself and its HP, and the performance of the obtained model on the MNIST’s validation set. The generated layers’ representation is obtained from the output layer of the RL’s actor, summarized in four values. These values represent the action space that is mapped first to a layer and then to that layer’s HPs. The model’s performance is represented with 32 values, each calculating the model’s accuracy results on each validation set’s batch of size 16. The heterogeneity in the dimensions of the action space and the performance requires a mapping to a uniform representation. This goal is achieved using a static NN that takes these inputs and produces a uniform output of 64 values, referred to as Intermediate Model Representation (IMR). When a layer and its performance are obtained, the state space, reflecting the current state of the environment, should also change. Since the state space represents a sequence of layers, the index corresponding to the generated layer is updated with the IMR. With more layers, each index is updated with a new representation. This way the state space reflects the two important pieces of information for each layer, its HPs, and its performance. This process was touched upon in Figure 1.
The stopping criteria for the RL agent involve either generating more than 6 layers, minimal improvement in the performance with the addition of layers (0.001), or accuracy below 60%. These criteria are defined to avoid any unnecessary generation of models. A Deep Deterministic Policy Gradient (DDPG) [14] handles the continuous action spaces in this environment. The critic and the actor consist of target and online subnets to avoid radical changes in critic and actor updates [14]. The training of RL agents involves exploration and exploitation of the gained knowledge. In DDPG, the agent explores by adding random noise to the sampled actions, so that the agent can experiment with different combinations of layers and obtain their performance results. To remove data correlations for the RL agent inherent in the sequential structure of the data, experience replay (ER) [14] buffer is utilized to store data during the exploration stage.
The training process is realized over several episodes. In each episode, 10 full models (i.e., until a stopping criterion is met) are generated via the utilization of CPU parallelization. Once the ER buffer is full, five RL optimization rounds are implemented in each episode. The number of episodes and the size of the ER buffer are important parameters for the HPO process. Their optimization is imperative to obtain models with good performances in a short period. During the HPO process, the training set consists of 20,000 images while the validation set includes 10,000 images out of the original 60,000 images. The experimental procedure is conducted on SHARCNET’s Graham cluster on a node with one V100 GPU, 64 Gbs of RAM, and 12 cores, such that each experiment does not exceed 7 days to avoid long queuing times. The implementation will be available on a GitHub repository upon publication.
V Results and Discussion
This section reports the results of the experimental procedure and analyzes them based on the defined performance metrics and layer-generation interpretability. TRL-HPO results are obtained in the exploitation phase of the offline RL.
V-A TRL-HPO vs. SOTA
Methods
Best Accuracy
Running Time
AccTime
TRL-HPO 99.1 99.3 94.6 % $\pm$ 3.8%
MetaQNN [5] 99.5 192-240 88.5 % $\pm$ 4.3%
Neary [8] 95.8 1.78 95.8%
TABLE II: TRL-HPO versus SOTA in terms of accuracy and running time
Table II shows the results of TRL-HPO versus SOTA methods that built CNN models from scratch. Based on evaluation results, MetaQNN produces the best model when classifying MNIST data reported in the exploration phase. This result is attributed to the freedom endowed to the controller to generate models up to 12 layers and the integration of priors via the addition of dropout layers every 2 layers [5]. The accumulation of more layers improves classification results; however, at the expense of wider search space that increases the running time. This is evident with the computational footprint of MetaQNN that exceeds TRL-HPO, despite running on 8$\times$ the amount of GPU resources. On the other hand, the approach in [8] is the first to converge based on the reported results, given it is a Multi-agent RL algorithm. However, this method experimented with four configurations, highlighting the limited exploration of this strategy and the pre-mature convergence of the master agent. This limitation is reflected by its poor best models compared to the other two approaches, excluding it from further analysis. Since the larger search space of MetaQNN would not provide a fair comparison with TRL-HPO, it was imperative to compare these two approaches using AccTime. Within TRL-HPO convergence time, its generated models exhibit the best performance compared to the SOTA models. Despite TRL-HPO’s need to generate more models given the requirement of progressive reward, it outperforms other models in the AccTime metric. Compared to other SOTA methods, TRL-HPO can progressively generate CNN models with satisfactory performance in a shorter period. These two conditions are important for resource-constrained IoT environments. In terms of AccTime, on average, TRL-HPO outperforms MetaQNN by 6.8 %.
V-B Layer Analysis
Investigating the effect of the addition of layers is central to understanding the contribution of each layer towards improving accuracy results and the combination of layers that yields the best performances. This analysis unveils the complex relationships between layers facilitating the transparency of the HPO process. Both goals were considered in the design of TRL-HPO, representing a key differentiating factor versus SOTA methods. To gain these insights, two questions need to be answered. (1) What are the layer combinations degrading the performance of CNN models? and (2) How are layer affinities reflected in the attention mechanism?
[…]
[scale=0.4]neg_reward_prev_layer.png
Figure 2: Layer affinity for models with negative rewards
Figure 2 summarizes the distribution of layer combinations with negative rewards. Two combinations stand out, Class 1 represented by Conv2D and Conv2D combinations and Class 2 represented by two consecutive FCLs. A layer combination refers to the layer that produced a negative reward and its previous layer, shedding light on the combinations that degrade model performance. Two main insights can be gathered from Figure 2. The first is that the negative reward is overwhelmingly attributed to the accumulation of FCLs. This means that the stacking of multiple of these layers is superfluous on the MNIST dataset, a conclusion that aligns with the best models obtained in [5]. In a similar vein, the second insight highlights that additional Conv2D layers can be detrimental to model performance, especially with MNIST data of limited visual complexity. As such, these observations are beneficial to garner knowledge about superfluous stacking of layers, suitable for environments with resource constraints.
[…]
[scale=0.4]diff_attn.png
Figure 3: Differences in attention between layers of best reward and layers with negative reward
Figure 3 depicts the distribution of the differences in attention values between layers with the best performances and layers with negative rewards. The distribution demonstrates the close fidelity of the attention mechanism and the utility of a specific layer in its classification performance. As such, the attention mechanism captures the close affinity between layers assigning higher weights to promising layers while shunning layers with poor performances. The difference in attention values averages 0.22, a significant value considering that some attention is assigned to the layers that are yet to be generated. In these cases, the attention values of the layer degrading model performances are higher than these to-be-generated layers. Therefore, we conclude that the attention mechanism reflects the relationship between layers within the context of generating best-performing models.
VI Open Challenges and Future Recommedations
The implementation of the TRL-HPO process opens new avenues for exploring the improvement of this process, which spans many research and practical questions about RL-based HPO implementations.
VI-A Computational Time
Despite improvements in the efficiency of the TRL-HPO process, it has yet to converge in a short time. The bottleneck in the training process originates from two main sources: (1) the training of individual models, and (2) optimization rounds of the actor-critic approach. The first source is a necessary evil because of the exploration requirements and the need to evaluate the actor’s actions. Determining the training data size, the number of epochs, and learning rates is required so that close to optimal accuracy is achieved. During the exploration and action assessment phases, repetitive models can be generated. Storing these models is a method to avoid retraining them, which can overwhelm the available GPU or RAM resources, especially with a large ER buffer. An alternative method is storing already trained models in a database or a hash table, whereby the model is represented via its HP or its hash value. A salient issue relates to querying algorithms that should search for HP representations available in the database. The querying time will increase with the expansion of the database, requiring a more intricate search procedure. These methods should be investigated to budget the computational time towards more fruitful procedures that benefit RL-based solutions.
VI-B Exploration
RL implementations depend on the trial-and-error procedure taking place in the explorations phase. However, a shortcoming of this phase, either through DDPG’s random noise or $\epsilon$-greedy algorithms [5], relates to its lack of direction. This means that the agent continues exploring unpromising areas of the search space with relatively poor performances. Therefore, to reduce this unnecessary exploration, it is imperative to include priors or rectify the exploration stage. As such, aspects of the BO process should be integrated into the RL-based strategies to streamline the exploration process. The randomness can be incorporated using a stochastic policy, which is part of the soft actor-critic policy gradient approach [15]. In the exploration stage, the stochastic policy outputs the mean and standard deviation of a normal distribution. The standard deviation is progressively reduced to achieve a deterministic policy based on the reward function. As such, this suggestion opens new frontiers toward examining methods that can associate the uncertainty of the stochastic policy with steering the RL’s agent exploration.
VI-C Reward Function
The progressive reward function devised in TRL-HPO formulation is foundational to gaining insights into the contributions of each layer to performance enhancement. However, this layer accumulation process exposes two limitations. The first limitation pertains to the training of many models, even with a single layer, to infer this performance. This pre-requisite restricts expanding the model generation to a few layers, which applies to our experimental procedure. These constraints can be sidestepped via block generation at the expense of sacrificing transparency and layer diversity. A block is defined as a set of similar basic layers. This small modification is reflected by altering the actor’s output to produce the number of these similar blocks. The second limitation of this process relates to the training procedure. The HPO procedure is conducted on a subset of the training set to facilitate the training and validation procedures. This means that the performance of the generated CNN models can significantly vary, depending on the train and test splits. In light of the stopping criteria, this factor is crucial in deciding whether to halt or continue the layer generation process.
The reward function that drives the model generation process can be adjusted to achieve better results. Instead of performance differences, the reward can be the inverse distance to a perfect metric (100% accuracy in our example). In this way, the rewards are less sparse but require the tuning of a discount rate. Additionally, other metrics, such as f1-score and top-3 accuracy can replace accumulated accuracy. The adoption of any of these metrics depends on the practitioners’ priorities.
VI-D Transferability
With each new contribution toward the HPO procedure, questions of generalizability and transferability loom to undermine their usefulness, which also applies to RL-based solutions. In TRL-HPO, the actor and critic are constructed to generate and evaluate models using basic layers of FCL, CNN, and MaxPool layers, necessary to build DNN models. The transferability comes into question when TRL-HPO trained on one dataset needs to be applied to other datasets. In that regard, the main concern relates to the size of the input sequence, trained on the initial dataset, which constrains the size of the produced models. This raises the question of how to expand the input sequence to involve more layers?.
The answer to this question requires re-defining the concept of the input sequence (state) supplied to the actor and critic. In its current form, each index in the state sequence combines the layer definition and its performance. To enable transferability and the expansion of TRL-HPO to more profound use cases, the actor should output an additional value representing the number of similar layers. The remaining alteration is to determine the upper bound of the permitted number of these layers so that the model’s capacity would match the dataset’s complexity.
VII Conclusion
The HPO process is a fundamental step in the ML pipeline that contributes to enhancing model performance. However, its computational footprint is prohibitive for widespread adoption and current methods overlook the transparency in the layer generation process. These factors are imperative to realize IoT applications’ function, AVs in particular, requiring model partitioning to fulfill the resource constraints of edge environments. To address these limitations, this paper proposes TRL-HPO framework that combines transformers with an RL actor-critic approach. The attention mechanism, parallelization, and the progressive generation of layers are all novel properties of this framework within the transparency and time requirements. These advantageous factors were empirically established compared to SOTA approaches using the MNIST dataset. Moreover, a list of research questions scrutinizing RL-based solutions, including TRL-HPO, is presented along with a set of recommendations. Future work will target the presented questions to design a general-purpose HPO process.

### v2

@@
 Convolutional Neural Networks (CNNs) are the staple implementation of neural networks (NNs) for image classification and object detection.
-The progress in this field is attributed to the increased complexity of CNN architectures in terms of the type of connections among different layers and their depth, which increased the computational demands.
-As a result, the accuracy of these models depends on the number and the type of layers, the connections between these layers, and the parameters assigned to each layer [1].
-In addition to accuracy, other important factors that dictate the design of CNNs include training time, inference time, and energy consumption, which are emphasized in the resource-constrained Internet of Things (IoT) environment.
-The number of possible choices for designing CNNs is extremely large, which is prohibitive for manual search.
-These circumstances have promoted the automated CNN architecture search, achieved through the field of Neural Architecture Search (NAS).
+The progress in this field is attributed to the increased complexity of CNN architectures in terms of the type of connections among different layers and their depth, which increased the computational demands [1].
+Therefore, the accuracy of these models depends on the number and the type of layers, the connections between these layers, and the parameters assigned to each layer [2].
+Other important factors dictating the design of CNNs include training time, inference time, and energy consumption, emphasized in the resource-constrained Internet of Things (IoT) environment.
+The number of possible choices for designing CNNs is extremely large, which promoted the automated CNN architecture search, achieved through Neural Architecture Search (NAS).
 The NAS field has recently seen great progress, due to the incorporation of Reinforcement Learning (RL) agents to search for the best CNN configurations.
-The RL’s appeal stems from its generalizability of hyper-parameter (HP) combinations via function approximation and the trial-and-error approach, which has the potential to reduce the computational demand of NAS [2].
-This field is dominated by the literature that either builds CNNs from scratch using basic NN layers or a pre-defined collection of these layers, or optimizes the HPs of an already-existing CNN.
+The RL’s appeal stems from its generalizability of hyper-parameter (HP) combinations via function approximation and the trial-and-error approach, which can reduce the computational demand of NAS [3].
+This field is dominated by the literature that either builds CNNs from scratch using basic NN layers or a pre-defined collection of these layers or optimizes the HPs of an already-existing CNN.
 Autonomous vehicles (AV) are part of the envisioned IoT applications that utilize edge servers of limited computational resources.
 The realization of AVs heavily depends on image classification models deployed on these servers.
-As such, the disparate computational capabilities of edge servers are prohibitive for the prolonged execution of large models.
+Their disparate computational capabilities are prohibitive for the prolonged execution of large models.
 Model partitioning is key to resolving these limitations toward fulfilling the promise of AVs [4].
@@
 In current implementations, this vision is stalled by multiple oversights.
-The sequential nature of model generation results from the dependence between layer combinations.
-This dependence can be reflected by the recurrent structures of Recurrent Neural Networks (RNNs).
+The sequential nature of model generation results from the dependence between layer combinations, which can be reflected by the recurrent structures of Recurrent Neural Networks (RNNs).
 Therefore, RNN-based controllers constructing CNN models dominated the RL-based NAS field [5, 6].
@@
 Within the constraints of computational resources and model partitioning requirements, the state-of-the-art (SOTA) approaches are limited by their impracticability and lack of transparency, hindering their deployment in real-world environments.
-This paper presents Transformer-based Reinforcement Learning Hyper-parameter Optimization (TRL-HPO) that alleviates the shortcomings of the current methods.
+This paper presents a Transformer-based Reinforcement Learning Hyper-parameter Optimization (TRL-HPO) model that alleviates the shortcomings of the current methods.
 The TRL-HPO addresses the transparency and long-computational times of RL-based solutions with competitive performances.
-The Transformer architecture overshadows the RNN-based methods by integrating the multi-headed self-attention mechanism [6].
+The Transformer architecture overshadows the RNN-based methods by integrating the multi-headed self-attention (MHSA) mechanism that enables parallelization [7], addressing the long computational times.
 The TRL-HPO controller equipped with the attention mechanism allows us to gain insights from the layer-generation process, enhancing its interpretability.
-Moreover, the transformers’ architecture supports parallelization, which addresses the long computation times of its RNN-based counterparts [6].
 The reward is produced with every generated layer to showcase improvement instead of waiting for termination conditions to unfold, a condition matching CNN partitioning requirements, as depicted in Figure 1.
 The contributions of this paper are as follows:
-  • Propose a novel Hyper-parameter Optimization (HPO) process named TRL-HPO that is the first to combine transformer architecture and actor-critic RL;
+  • Propose a novel Hyper-parameter Optimization (HPO) process named TRL-HPO that is the first to combine transformer architecture and actor-critic (AC) RL;
   • Enhance the CNN’s model performance generated within a shorter period compared to SOTA approaches using the TRL-HPO process;
@@
 Bayesian Optimization (BO), Evolutionary Search (ES) algorithms, and RL agents are the three main tools for HPO implementation.
-BO and ES methods are limited by their assumptions [1] and lack of generality [7], which favors the application of RL techniques.
+BO and ES methods are limited by their assumptions [2] and lack of generality [8], which favors RL techniques.
 The works of Baker et al.
@@
 MONAS [11] considers the energy and accuracy constraints, whereas MNASNet [12] integrates the inference latency of developed models.
-The shortcomings exhibited by the integration of RL into the HPO process in these implementations hinder their deployment in IoT environments.
+The shortcomings of RL integration into the HPO process hinder their deployment in IoT environments.
 First, the adopted models lack the transparency that shows the interdependence between different layers.
-Second, some of these approaches take considerable time to converge, which is a privilege that is not always available in resource-constrained environments.
+Second, the long convergence times are prohibitive for deploying these models in resource-constrained environments.
 Third, some works focus on optimizing the HPs of a specific layer, reducing the state space at the expense of the layer’s diversity.
-Lastly, many of the works include prior knowledge instilled into the layer-generation process, such as the addition of dropout layers [4, 5].
-These limitations are addressed in this paper by proposing a transformer-based RL controller and the reward function formulation.
-The multi-headed attention mechanism facilitates the training process and adds transparency to the model generation process by analyzing the attention values.
-On the other hand, the reward function reflects the contribution of each layer to enhancing the classification results, a condition favorable to CNN model partitioning.
-[…]
-[scale=0.45]architecture_v3.PNG
+Lastly, many works include prior knowledge in the layer-generation process, such as the addition of dropout layers [5, 6].
+This paper addresses these limitations by proposing a transformer-based RL controller and the reward function formulation.
+The MHSA facilitates the training process and adds transparency to the model generation by analyzing the attention values.
+On the other hand, the reward function reflects the contribution of each layer to enhancing the classification results, which favors CNN model partitioning.
 Figure 1: TRL-HPO framework
 III Methodology: TRL-HPO
-This section explains the building blocks of TRL-HPO, including the transformer and RL’s actor-critic architectures, which are shown in Figure 1, and the motivation for this combination toward a more efficient and transparent HPO process.
+This section explains the building blocks of TRL-HPO, including the transformer and RL’s AC architectures, which are shown in Figure 1, and the motivation for this combination toward a more efficient and transparent HPO process.
 III-A Transformer
@@
 This architecture enables parallelization accelerating the training of transformer models.
-The building blocks of transformers include attention mechanism, multi-headed attention, position-wise feed-forward network, and positional encoding (PE).
-The transformers follow the encoder-decoder structure thereby each of these structures is formed by stacking combinations of these identical layers.
-The ordering of these layers with respect to the encoder and decoder is depicted in Figure 1.
-The Positional Encoding computes the sequence’s order and is added to the input embedding of the encoder and decoder stacks so that the order is incorporated into the input and output data.
-To obtain a unique order, a sinusoidal function takes as inputs the position and the embedding dimensions.
-This structure extrapolates to sequences that are longer than the ones encountered during training [6].
-The self-attention module calculates the weights assigned by each input sequence to other sequences.
-This way the recurrence structure is avoided because self-attention reflects the dependencies between input and specific output and improves the modelling of long-range dependencies [6].
-As such, the attention mechanism resembles a fully connected layer (FCL), such that the weights reflect the pairwise relationship from previous inputs.
-To fully exploit this method, the transformer includes multi-headed attention that replicates the attention calculation.
-This process generates different representations of the input data at different positions.
-The calculation of these attention matrices is independent and facilitates their parallelization.
-Lastly, a feed-forward neural network is used as the subsequent layers to the multi-head attention mechanism.
+The building blocks of transformers include attention mechanism, multi-headed attention (MHSA), position-wise feed-forward network, and positional encoding (PE).
+The transformers follow the encoder-decoder structure thereby each of these structures is formed by stacking combinations of these identical layers, as depicted in Figure 1.
+The PE computes the sequence’s order and is added to the input embedding of the encoder and decoder stacks so that the order is incorporated into the input and output data.
+To obtain a unique order, a sinusoidal function inputs the position and the embedding dimensions.
+The self-attention module calculates the weights each input sequence assigns to other sequences.
+This way the self-attention reflects the dependencies between input and specific output and improves the modelling of long-range dependencies [7].
+As such, the attention mechanism resembles a fully connected layer (FCL), whereby the weights reflect the pairwise relationship from previous inputs.
+To fully exploit this method, the transformer includes MHSA that splits the attention calculation among different heads of the embedding that can be calculated independently, facilitating its parallelization.
+Lastly, a feed-forward neural network (FFNN) is the subsequent layer to the MHSA.
 III-B Actor-Critic Reinforcement Learning
@@
 Its actions are refined based on a reward function that signals the utility of these actions [13].
-In connection to the HPO process, the RL’s experimentation with the absence of any prior knowledge mirrors the inability to hypothesize the best combination of basic CNN layers.
-The only knowledge for the HPO process relates to the order of the stacked layer that starts with grid-like inputs and ends with a fully connected layer (FCL).
-As such, the combination of unknown orientation and trial-and-error experimentation matches the requirements of the HPO process.
+The RL’s experimentation with the absence of prior knowledge mirrors the ignorance of the best combination of basic CNN layers.
+The only knowledge for the HPO process relates to the order of the stacked layer that starts with grid-like inputs and ends with an FCL.
+The combination of unknown orientation and trial-and-error experimentation matches the requirements of the HPO process.
 The RL methods can be split into value-based and policy-based methods.
@@
 The high variability of value estimations and their under-performance in scenarios with continuous action spaces promotes policy-based methods [13].
-The policy-based methods focus on optimizing the policy to maximize the expected cumulative rewards [12].
+The policy-based methods optimize the policy to maximize the expected cumulative rewards [13].
 However, the latter methods suffer from sample inefficiency, which is an asset of value-based methods.
-To complement these methods’ advantages, an actor-critic approach combines both policy-based (actor) and value-based (critic) methods [12].
+To complement these methods’ advantages, an actor-critic (AC) approach combines both policy-based (actor) and value-based (critic) methods [13].
 Projected to the HPO problem, the actor outputs an action that maps to a layer and its HPs.
-The critic evaluates this action by outputting a value showing the quality of the taken action [12].
+The critic evaluates this action by outputting a value showing the action’s quality [13].
 III-C Transformers and Actor-critic RL
-As the name suggests, the proposed framework, Transformer-based Reinforcement Learning HPO (TRL-HPO) is the convergence of transformers and actor-critic RL.
-The actor and the critic are implemented using the transformer architecture.
+The proposed framework, Transformer-based Reinforcement Learning HPO (TRL-HPO) is the convergence of transformers and AC RL.
+The actor and the critic are implemented using the transformer’s decoder architecture.
 The steps involved in the actions’ generation, CNN model construction, and their evaluation are depicted in Figure 1.
-The combination of TRL-HPO enables the RL agent to harness the strengths of transformers to the benefit of the HPO process.
+The combination of TRL-HPO enables the RL agent to harness the strengths of transformers to benefit the HPO process.
 This new framework is the first trial to integrate transformers into the HPO process, opening a new frontier toward the exploitation of this novel architecture to address a lingering problem in the field of ML.
-This experiment does not plan to only exhibit the viability of transformers’ implementation to HPO, but it rather analyzes the suitability of this architecture to the HPO process.
-We analyze TRL-HPO based on its inherent advantages and its benefits versus the SOTA approaches.
-In terms of the framework, two main advantages are reaped.
+This experiment evaluates the viability and suitability of this architecture to the HPO process.
+We analyze TRL-HPO based on its inherent advantages and benefits versus the SOTA approaches.
+Two main advantages are reaped.
 The first is that TRL-HPO builds models from scratch, rendering it a general-purpose method.
-The proof of concept is constrained to a single use case; however, the framework can be expanded to any problem that utilizes DNN models.
+The proof of concept is constrained to a single use case; however, the framework can be expanded to any DNN-related problem.
 The second concerns the definition of the reward function that is progressively updated with each stacked layer.
@@
 This feature fits the requirements of environments that seek to balance resource constraints and accuracy objectives.
-The transformer architecture results in two inherent advantages to the HPO process.
-The first is connected to the parallelization enabled by the multi-headed attention mechanism, which reduces the running time of the HPO process.
-The second benefit stems from the attention mechanism that showcases the relationships between generated layers.
-A by-product of this feature is the inference of the combination of layers that produce better results, which enhances the transparency of the HPO process.
+The transformer architecture results in two main advantages to the HPO process.
+The first relates to handling long sequences, which benefits the transferability of TRL-HPO from small to larger datasets.
+The second benefit relates to the MHSA, which facilitates the parallelization of TRL-HPO and reduces its running time.
+Additionally, MHSA reveals the relationships between generated layers, which allows the inference of the combination of layers that produce better results, enhancing the transparency of the HPO process.
 IV Implementation Details
-The evaluation of TRL-HPO is conducted on the MNIST dataset [13].
-This dataset is a large database of handwritten digits, from 0 to 9, containing 60,000 training images and 10,000 testing images.
+The evaluation of TRL-HPO is conducted on the MNIST dataset [14], which is a large database of handwritten digits, from 0 to 9, containing 60,000 training and 10,000 testing images.
 Each image is a greyscale image of 28 $\times$ 28 dimensions.
-This dataset serves as a good proof-of-concept for the proposed framework, due to its availability and limited complexity.
+The availability and limited complexity of MNIST can serve as a good proof-of-concept for TRL-HPO.
 The former factor enables a fair comparison with other SOTA approaches, such as [6, 9].
@@
 The comparison with SOTA approaches is based on the convergence time to the best solution (hrs), the classification accuracy (%), and the best CNN classification model (%) obtained by each SOTA approach upon the completion of TRL-HPO process (AccTime).
-We compare TRL-HPO with Baker et al.
-[5] and Neary [8], which are the only RL-based solutions that build CNN models from scratch.
+Reproducibility is a major issue that plagues the HPO process [2].
+With this factor in mind, the choice of competing methods depended on two conditions: (1) The availability of source code that enables a fair comparison with TRL-HPO and dispels any introduced biases by implementing a method from scratch.
+(2) The similarity in the experimental procedure for building CNNs from scratch using an RL-based method and the availability of results applied to the MNIST dataset.
+Therefore, we compare TRL-HPO with Baker et al.
+[6] and Neary [9], which abide by at least one of these conditions.
 Layer HP Values
@@
 Stacking a combination of these layers adapts to the input structure.
-In connection to the RL agent, the design of the state and action space and the reward function is imperative to drive the RL agents’ action generation, which requires defining the transformers’ input sequence.
+The state and action space design and the reward function are imperative to drive the RL agents’ action generation, which requires defining the transformers’ input sequence.
 The reward function represents a layer’s contribution towards improving the classification results, which means that the DNN model’s first layer produces the highest rewards compared to any subsequent layers.
@@
 The generated layers’ representation is obtained from the output layer of the RL’s actor, summarized in four values.
-These values represent the action space that is mapped first to a layer and then to that layer’s HPs.
+These values represent the action space mapped first to a layer and then to that layer’s HPs.
 The model’s performance is represented with 32 values, each calculating the model’s accuracy results on each validation set’s batch of size 16.
@@
 This process was touched upon in Figure 1.
-The stopping criteria for the RL agent involve either generating more than 6 layers, minimal improvement in the performance with the addition of layers (0.001), or accuracy below 60%.
+The stopping criteria for the RL agent involve generating more than 6 layers, minimal improvement in the performance with the addition of layers (0.001), or accuracy below 60%.
 These criteria are defined to avoid any unnecessary generation of models.
@@
 The critic and the actor consist of target and online subnets to avoid radical changes in critic and actor updates [13].
-The training of RL agents involves exploration and exploitation of the gained knowledge.
 In DDPG, the agent explores by adding random noise to the sampled actions, so that the agent can experiment with different combinations of layers and obtain their performance results.
 To remove data correlations for the RL agent inherent in the sequential structure of the data, experience replay (ER) [13] buffer is utilized to store data during the exploration stage.
+The actor and the critic follow the transformer architecture, depicted in Figure 1.
+Two main differences highlight these components’ distinctive roles: the transformer’s output and the learning rate (lr).
+The actor’s output is in the [0, 1] range mapped to a layer, while the critic’s is in the [-1, 1] range representing the state’s Q-value.
+Regarding the lr, it is recommended that the actors’ (1e-5) be slower than the critics’ (1e-4) [13].
+The input embedding’s dimension for each input space is equal to 64.
+The embeddings are inputs to multiple encoding layers, equivalent to 2 in our implementation, each constituting a Transformer block, represented by the MHSA and FFNN.
+Each of these blocks has the same number of input and output dimensions.
+The definition of MHSA and FFNN requires highlighting two values, number of heads equal to 4 and expansion factor equal to 4.
 The training process is realized over several episodes.
-In each episode, 10 full models (i.e., until a stopping criterion is met) are generated via the utilization of CPU parallelization.
+In each episode, 10 full models are generated via CPU parallelization.
 Once the ER buffer is full, five RL optimization rounds are implemented in each episode.
 The number of episodes and the size of the ER buffer are important parameters for the HPO process.
-Their optimization is imperative to obtain models with good performances in a short period.
+Their optimization is imperative to obtain models with good performances in a short time.
 During the HPO process, the training set consists of 20,000 images while the validation set includes 10,000 images out of the original 60,000 images.
 The experimental procedure is conducted on SHARCNET’s Graham cluster on a node with one V100 GPU, 64 Gbs of RAM, and 12 cores, such that each experiment does not exceed 7 days to avoid long queuing times.
-The implementation will be available on a GitHub repository upon publication.
+The implementation is available on the GitHub repository [^1] https://github.com/Western-OC2-Lab/TRL-HPO.
 V Results and Discussion
-This section reports the results of the experimental procedure and analyzes them based on the defined performance metrics and layer-generation interpretability.
-TRL-HPO results are obtained in the exploitation phase of the offline RL.
+This section reports and analyzes the results based on the defined performance metrics and layer-generation interpretability. TRL-HPO results are obtained in the exploitation phase of the offline RL.
 V-A TRL-HPO vs. SOTA
@@
 Based on evaluation results, MetaQNN produces the best model when classifying MNIST data reported in the exploration phase.
-This result is attributed to the freedom endowed to the controller to generate models up to 12 layers and the integration of priors via the addition of dropout layers every 2 layers [5].
+This result is due to the controller’s freedom to generate models up to 12 layers and the integration of priors by adding dropout layers [6].
 The accumulation of more layers improves classification results; however, at the expense of wider search space that increases the running time.
 This is evident with the computational footprint of MetaQNN that exceeds TRL-HPO, despite running on 8$\times$ the amount of GPU resources.
-On the other hand, the approach in [8] is the first to converge based on the reported results, given it is a Multi-agent RL algorithm.
+On the other hand, the approach in [9] is the first to converge based on the reported results.
 However, this method experimented with four configurations, highlighting the limited exploration of this strategy and the pre-mature convergence of the master agent.
@@
 These two conditions are important for resource-constrained IoT environments.
-In terms of AccTime, on average, TRL-HPO outperforms MetaQNN by 6.8 %.
+In terms of AccTime, on average, TRL-HPO outperforms MetaQNN by 6.8 % in accuracy.
+The improvement in the accuracy result of TRL-HPO compared to MetaQNN is statistically significant (p-value = 0.02) by applying a t-test on the distribution of TRL-HPO versus the 1-sd values of MetaQNN.
+In a resource-constrained IoT environment, analyzing the energy consumption and processing power of TRL-HPO compared to SOTA methods is imperative.
+We analyze these methods using the method’s number of parameters and the floating point operations (FLOPs), which cover IoT-related concerns.
+The MetaQNN requires storing all models and state transitions to find the best-performing models, which with many combinations of layers and HPs, is prohibitive for the IoT environment.
+In Neary et al.
+approach [9], an RL agent is assigned to every HPs and a master agent finds the best combination of individual HPs.
+For Neary et al., 399K FLOPs are required for inference whereas TRL-HPO requires 991k FLOPs.
+On the other hand, both models have the same number of model parameters (78.8k parameters).
 V-B Layer Analysis
@@
 and (2) How are layer affinities reflected in the attention mechanism?
-[…]
-[scale=0.4]neg_reward_prev_layer.png
 Figure 2: Layer affinity for models with negative rewards
@@
 Two combinations stand out, Class 1 represented by Conv2D and Conv2D combinations and Class 2 represented by two consecutive FCLs.
-A layer combination refers to the layer that produced a negative reward and its previous layer, shedding light on the combinations that degrade model performance.
+A layer combination refers to the layer that produced a negative reward and its previous layer.
 Two main insights can be gathered from Figure 2.
 The first is that the negative reward is overwhelmingly attributed to the accumulation of FCLs.
-This means that the stacking of multiple of these layers is superfluous on the MNIST dataset, a conclusion that aligns with the best models obtained in [5].
-In a similar vein, the second insight highlights that additional Conv2D layers can be detrimental to model performance, especially with MNIST data of limited visual complexity.
-As such, these observations are beneficial to garner knowledge about superfluous stacking of layers, suitable for environments with resource constraints.
-[…]
-[scale=0.4]diff_attn.png
+This means that the stacking of multiple FCL layers is superfluous on the MNIST dataset, a conclusion that aligns with the best models obtained in [6].
+The second insight highlights that additional Conv2D layers can degrade model performance, especially with MNIST data of limited visual complexity.
+As such, these observations are beneficial to garner knowledge about stacking of layers, suitable for environments with resource constraints.
 Figure 3: Differences in attention between layers of best reward and layers with negative reward
@@
 As such, the attention mechanism captures the close affinity between layers assigning higher weights to promising layers while shunning layers with poor performances.
-The difference in attention values averages 0.22, a significant value considering that some attention is assigned to the layers that are yet to be generated.
+The difference in attention values averages 0.22, a significant value considering that some attention is assigned to the layers yet to be generated.
 In these cases, the attention values of the layer degrading model performances are higher than these to-be-generated layers.
 Therefore, we conclude that the attention mechanism reflects the relationship between layers within the context of generating best-performing models.
-VI Open Challenges and Future Recommedations
-The implementation of the TRL-HPO process opens new avenues for exploring the improvement of this process, which spans many research and practical questions about RL-based HPO implementations.
+Challenge Summary Solutions
+Computational Time Training of individual models Optimization rounds of actor-critic approach Storing the trained models in a hash table to avoid re-training
+Exploration Strategy DDPG’s lack of direction Application of RL methods with inherent randomness such as soft actor-critic approach
+Reward Function Training many models to infer their contribution to classification results Prioritization of accuracy Generating a block of similar layers Inclusion of inference time, energy consumption, and number of operations per second
+Reward Function Training many models to infer their contribution to classification results Prioritization of accuracy Generating a block of similar layers Inclusion of inference time, energy consumption, and number of operations per second
+Generalizability, Scalability, and Transferability Scalability and Transferability concerns with larger datasets and complex architectures.
+Generalizability to different IoT Environments Expansion of the input sequence and actor’s block generation Changing the mappings of the actor’s output to adapt to different DNN models
+Integration into Automated ML N/A Replacing the HPO process for IoT deployment frameworks
+TABLE III: Summary of Challenges
+VI Open Challenges and Future Recommendations
+The implementation of the TRL-HPO process opens new avenues for exploring the improvement of this process, which spans many research and practical questions about RL-based HPO implementations, summarized in Table III.
 VI-A Computational Time
 Despite improvements in the efficiency of the TRL-HPO process, it has yet to converge in a short time.
-The bottleneck in the training process originates from two main sources: (1) the training of individual models, and (2) optimization rounds of the actor-critic approach.
+The bottleneck in the training process originates from two main sources: (1) the training of individual models, and (2) optimization rounds of the RL approach.
 The first source is a necessary evil because of the exploration requirements and the need to evaluate the actor’s actions.
-Determining the training data size, the number of epochs, and learning rates is required so that close to optimal accuracy is achieved.
+Determining the training data size, the number of epochs, and learning rates are required to achieve close to optimal accuracy.
 During the exploration and action assessment phases, repetitive models can be generated.
-Storing these models is a method to avoid retraining them, which can overwhelm the available GPU or RAM resources, especially with a large ER buffer.
+Storing these models can avoid retraining them, which can overwhelm the available GPU or RAM resources, especially with a large ER buffer.
 An alternative method is storing already trained models in a database or a hash table, whereby the model is represented via its HP or its hash value.
@@
 VI-B Exploration
-RL implementations depend on the trial-and-error procedure taking place in the explorations phase.
+RL implementations depend on the trial-and-error procedure in the exploration phase.
 However, a shortcoming of this phase, either through DDPG’s random noise or $\epsilon$-greedy algorithms [6], relates to its lack of direction.
 This means that the agent continues exploring unpromising areas of the search space with relatively poor performances.
-Therefore, to reduce this unnecessary exploration, it is imperative to include priors or rectify the exploration stage.
-As such, aspects of the BO process should be integrated into the RL-based strategies to streamline the exploration process.
+Therefore, it is imperative to include priors or rectify the exploration stage to reduce unnecessary exploration.
+Aspects of the BO process should be integrated into the RL-based strategies to streamline the exploration process.
 The randomness can be incorporated using a stochastic policy, which is part of the soft actor-critic policy gradient approach [15].
-In the exploration stage, the stochastic policy outputs the mean and standard deviation of a normal distribution.
+In the exploration stage, the stochastic policy outputs the normal distribution’s mean and standard deviation.
 The standard deviation is progressively reduced to achieve a deterministic policy based on the reward function.
-As such, this suggestion opens new frontiers toward examining methods that can associate the uncertainty of the stochastic policy with steering the RL’s agent exploration.
+This suggestion opens new frontiers toward examining methods that can associate the uncertainty of the stochastic policy with steering the RL’s agent exploration.
 VI-C Reward Function
-The progressive reward function devised in TRL-HPO formulation is foundational to gaining insights into the contributions of each layer to performance enhancement.
-However, this layer accumulation process exposes two limitations.
-The first limitation pertains to the training of many models, even with a single layer, to infer this performance.
-This pre-requisite restricts expanding the model generation to a few layers, which applies to our experimental procedure.
-These constraints can be sidestepped via block generation at the expense of sacrificing transparency and layer diversity.
-A block is defined as a set of similar basic layers.
-This small modification is reflected by altering the actor’s output to produce the number of these similar blocks.
-The second limitation of this process relates to the training procedure.
-The HPO procedure is conducted on a subset of the training set to facilitate the training and validation procedures.
-This means that the performance of the generated CNN models can significantly vary, depending on the train and test splits.
-In light of the stopping criteria, this factor is crucial in deciding whether to halt or continue the layer generation process.
-The reward function that drives the model generation process can be adjusted to achieve better results.
-Instead of performance differences, the reward can be the inverse distance to a perfect metric (100% accuracy in our example).
-In this way, the rewards are less sparse but require the tuning of a discount rate.
-Additionally, other metrics, such as f1-score and top-3 accuracy can replace accumulated accuracy.
-The adoption of any of these metrics depends on the practitioners’ priorities.
-VI-D Transferability
-With each new contribution toward the HPO procedure, questions of generalizability and transferability loom to undermine their usefulness, which also applies to RL-based solutions.
-In TRL-HPO, the actor and critic are constructed to generate and evaluate models using basic layers of FCL, CNN, and MaxPool layers, necessary to build DNN models.
-The transferability comes into question when TRL-HPO trained on one dataset needs to be applied to other datasets.
-In that regard, the main concern relates to the size of the input sequence, trained on the initial dataset, which constrains the size of the produced models.
-This raises the question of how to expand the input sequence to involve more layers?.
-The answer to this question requires re-defining the concept of the input sequence (state) supplied to the actor and critic.
-In its current form, each index in the state sequence combines the layer definition and its performance.
-To enable transferability and the expansion of TRL-HPO to more profound use cases, the actor should output an additional value representing the number of similar layers.
-The remaining alteration is to determine the upper bound of the permitted number of these layers so that the model’s capacity would match the dataset’s complexity.
+The progressive reward function of TRL-HPO is foundational to gaining insights into the contribution of each layer to performance enhancement.
+However, this layer accumulation process requires training many models to infer their performance, restricting the model generation process to a few layers.
+These constraints can be sidestepped by generating similar layers at once, which is referred to as a block generation procedure.
+As such, computational time is gained at the expense of transparency and layer diversity.
+In resource-constrained environments, factors such as energy consumption, inference time, and number of computational operations per second can gain precedence over accuracy.
+These factors can be added to the reward function and assigned a weight based on application requirements.
+VI-D Generalizability, Scalability, and Transferability
+With each contribution toward the HPO procedure, questions of generalizability, scalability and transferability loom to undermine their usefulness; a concern that applies to RL-based solutions.
+In TRL-HPO, the actor and the critic generate and evaluate models using FCL, CNN, and MaxPool layers, necessary to build DNN models.
+When applied to complex datasets and architectures, skip and residual connections can be integrated into the TRL-HPO architecture.
+Towards that end, two main approaches can be followed: (1) expand the input space to include more layers, which is reflected in the sequence length variable, or (2) change the definition of the actor’s output from the layer and its HPs to include the number of similar layers and their HPs.
+Adopting any of these approaches should consider the trade-off between the availability of computational resources and the layer generation process’s transparency.
+The same adaptation procedure can be followed when confronting a transferability concern.
+The IoT environment is integrated into various fields, including smart cities, healthcare, buildings, and electric grids.
+This environment suffers from computational and communication resource scarcity, which can be detrimental to ML applications, such as anomaly detection, object detection, and forecasting.
+In connection to the IoT environment limitations, the TRL-HPO framework offers solutions manifested in three aspects: (1) its generalizability to different types of DNN models, including CNN, Long short-term memory (LSTM) and FFNN by changing the mappings of the actor’s output, (2) its low computational, storage, and processing footprint, compared to SOTA approaches, and (3) its transparency that enables the reliance on part and not all of the CNN models and exchanging model parameters when the need arises.
+VI-E Integration into AutoML
+The HPO process is an important procedure in the AutoML pipeline and its enhancement is central to the wide-scale deployment and adoption of these pipelines.
+The TRL-HPO process is envisioned to replace the generic HPO processes that are part of the ubiquitous cloud computing ML deployment modules.
+These modules enhance the efficiency and accessibility of ML deployment, by providing user-friendly deployment strategies.
+TRL-HPO that outperformed SOTA approaches in its convergence time while maintaining good accuracy results benefits experts who have budgetary constraints and desire to prove the viability of their products/models for customers/investors alike.
 VII Conclusion
-The HPO process is a fundamental step in the ML pipeline that contributes to enhancing model performance.
+The HPO process is a fundamental step in the ML pipeline that enhances model performance.
 However, its computational footprint is prohibitive for widespread adoption and current methods overlook the transparency in the layer generation process.
@@
 These advantageous factors were empirically established compared to SOTA approaches using the MNIST dataset.
-Moreover, a list of research questions scrutinizing RL-based solutions, including TRL-HPO, is presented along with a set of recommendations.
+Moreover, a list of research questions scrutinizing RL-based solutions, including TRL-HPO, is presented with corresponding recommendations.
 Future work will target the presented questions to design a general-purpose HPO process.
+VIII Acknowledgement
+This work was made possible by the facilities of the Shared Hierarchical Academic Research Computing Network (SHARCNET:www.sharcnet.ca) and Compute/Calcul Canada.